# Self-induced stochastic resonance: A physics-informed machine learning approach

### v1

@@ 5.4 Model, optimizer setup, and hyperparameters
Training is performed via backpropagation using the Adam optimizer, see the algorithm in the Appendix Appendix. We initialize a StatePredictor with three 128-unit Tanh blocks and a dynamics head and set $dt=0.05$, epochs $1\times 10^{4}$, batch size =512=512, and learning rate $1.0\times 10^{-3}$. The total loss is a weighted sum of the data loss, the initial-condition loss, the residual loss, and the barrier-based physics loss weighted by, $\lambda_{\tiny{data}},\lambda_{\tiny{ic}},\lambda_{\tiny{phy1}},\lambda_{\tiny{phy2}}$ respectively. During training, each minibatch is a consecutive time sequence.
For evaluation, we predict the target sequence and compute the normalized root-mean-square error (NRMSE). To ensure reproducibility, the NumPy and PyTorch random seeds are fixed to 42. The optimal training epoch is selected as the one yielding the smallest NRMSE, defined for a reference sequence $y=\{y_{t}\}_{t=1}^{N}$ and its predictions $\hat{y}=\{\hat{y}_{t}\}_{t=1}^{N}$ by
@@
where lower values indicate better predictive performance.
@@ 5.5 Results on training and evaluation
We evaluate the NASP–PINN on held-out sequences to measure both one-step prediction quality and multi-step rollout fidelity, and we perform an ablation over the loss components to quantify the contribution of the physics terms. Unless stated otherwise, training uses consecutive windows from simulated trajectories with parameters given in Sections 5.35.4; performance is reported as normalized root-mean-square error (NRMSE, lower is better), cf. Eq. (22). The evaluation phase includes both teacher-forced (one-step) prediction and open-loop (free-rollout) simulation, which together assess short-term fit and long-horizon stability.
We benchmark four NASP–PINN variants to quantify how progressively injecting physics affects learning dynamics and out-of-sample generalization. Figure 8 shows the NRMSE trajectory over epochs for four loss configurations: data only (red), data+ic+Phy1 (yellow), data+Phy2 (blue), and data+ic+Phy1+Phy2 (green). Adding physics consistently accelerates optimization and lowers the asymptotic error relative to a purely data-driven baseline. The full PINN (all terms) converges fastest and to the lowest error.
@@
Figure 8: Training performance of NASP–PINN under different loss configurations. Shown are NRMSE curves for four variants: data only (red), data+ic+Phy1 (yellow), data+Phy2 (blue), and data+ic+Phy1+Phy2 (green). Physics-informed terms (Phy1, Phy2) accelerate convergence and reduce error, with the full model achieving fastest convergence and the smallest generalization gap, and most stable training.
@@
Table 1 summarizes train/test NRMSE for the same variants. Data-only training yields the weakest generalization (0.053 train vs. 0.062 test), indicating mild overfitting to short-horizon local patterns. Enforcing dynamical residuals (Phy1) and/or barrier-matching (Phy2) narrows the generalization gap and lowers both errors (0.040–0.042 train, 0.050 test). The full loss achieves the best results by a clear margin (0.027 train, 0.035 test), demonstrating that the two physics priors are complementary: Phy1 aligns local derivatives with the SDE, while Phy2 shapes global exit statistics via timescale matching, together constraining the hypothesis class to trajectories that are both locally consistent and globally coherent. Thus, relative to data-only training, the PINN reduces test NRMSE by approximately 43% (0.062 $\rightarrow$ 0.035) while also lowering train error, consistent with a better-conditioned objective and stronger inductive bias. This quantitative improvement highlights how physics-informed constraints serve as effective regularizers that enhance both convergence speed and out-of-sample generalization.
@@
Table 1: NRMSE for NASP variants under different loss combinations
@@
From a learning-theoretic perspective, the physics terms act as inductive biases and implicit regularizers. Phy1 reduces variance by penalizing derivative mismatches estimated via automatic differentiation; Phy2 reduces bias in the noise-triggered event timing by anchoring the learned dynamics to the Arrhenius-type scaling required for SISR (timescale matching). Their combination explains the lower test error and the smoother, faster descent curves in Fig. 8. In broader machine-learning terms, these physics-based penalties constrain the hypothesis space in a way analogous to structural regularization in deep sequence models, improving both data efficiency and temporal consistency.
To test stability beyond the training horizon, we perform open-loop (free-rollout) simulations and compare against direct stochastic simulations. Figure 9 reports time series of $v(t)$ and $w(t)$ across training and prediction phases. The full PINN tracks both the fast spikes in $v(t)$ and the slow drift in $w(t)$ with close agreement, maintaining phase alignment and amplitude over an extended horizon. This indicates that the learned integrator is noise-aware and stiffness-compatible, two properties that are critical for slow–fast SDEs. The qualitative match of spike timing and recovery excursions corroborates that barrier-matching (Phy2) has taught the model where exits occur in $w$ and at which rate, while the residual loss (Phy1) controls local consistency between steps.
Figure 9: Rollout performance of NASP–PINN on test trajectories. Predicted and true SISR time series of $v(t)$ and $w(t)$ are shown for both training and open-loop phases. The full model (data+ic+Phy1+Phy2) accurately reproduces the true SISR dynamics, matching spike timing and recovery patterns while maintaining phase and amplitude over long horizons.
@@
Overall, the findings demonstrate that embedding the governing stochastic dynamics and SISR-specific constraints directly into the loss function transforms NASP into a data-efficient and generalizable framework for modeling noise-driven multiscale dynamics. Across all evaluation metrics—including learning curves, final NRMSE, and long-horizon rollouts—the Data + ic + Phy1 + Phy2 configuration consistently yields superior performance, with the most pronounced improvements observed in extended predictions where preserving the intrinsic slow–fast structure and accurate escape timing is critical for reproducing SISR. Physics-informed training, particularly the combination of local residual constraints and global barrier-matching, thus serves as an effective inductive bias that enhances generalization while ensuring the model faithfully respects the physics that drives self-induced stochastic resonance.
Figure 10 compares the noise–coherence relationship predicted by the NASP PINN with that obtained from direct stochastic simulations of the FHN model Eq. (1). The $\mathrm{CV}$ of interspike intervals is plotted as a function of noise intensity $\sigma$ for different values of the excitability parameter $a$ and the timescale separation $\varepsilon$. Panels (a1)–(a2) display the true $\mathrm{CV}$ curves computed from numerical simulations, whereas panels (b1)–(b2) show the corresponding NASP PINN-predicted results. The close quantitative agreement between both sets of curves demonstrates that the PINN accurately learns the underlying stochastic dynamics and correctly captures how the coherence of SISR depends on excitability and timescale separation. In particular, the network reproduces the characteristic shift of the $\mathrm{CV}$ minima toward higher noise intensities as $a$ or $\varepsilon$ increase, reflecting the loss of timescale matching between deterministic relaxation and stochastic escape predicted by the theoretical SISR condition. This confirms that embedding the governing physics directly into the training objective enables the PINN to generalize beyond the training data and faithfully recover the resonance structure of the stochastic FHN system.
@@
Figure 10: True and NASP PINN-predicted coherence of SISR. Panels (a1)–(a2) show $\mathrm{CV}$ versus noise intensity $\sigma$ for varying excitability ($a$) and timescale separation ($\varepsilon$) obtained from direct stochastic simulations, while (b1)–(b2) show the corresponding PINN predictions. The PINN faithfully reproduces the noise–coherence relationship of the FHN system, demonstrating its ability to capture the dependence of SISR on both excitability and timescale separation. Parameters: $b=1.0$, $c=2.0$.
@@ 6 Summary and conclusions
Building on these theoretical insights, we introduced a Physics-Informed Neural Network (PINN) based on a Noise-Augmented State Predictor (NASP) architecture to learn the stochastic FHN dynamics directly from data. The composite loss functional integrated data fidelity, dynamical residuals, and physics-based constraints derived from the SISR conditions. Training results demonstrated that incorporating both local dynamical consistency and global timescale-matching constraints significantly improved accuracy, convergence speed, and generalization compared with purely data-driven approaches. The trained PINN accurately reproduced long-term SISR dynamics and captured the dependence of the CV on noise intensity, excitability, and timescale separation, validating its ability to encode the underlying stochastic physics.
@@
In addition, this physics-informed neural framework aligns with current trends in scientific machine learning, where neural ordinary differential equations (neural ODEs) [56], reservoir computing [57], and physics-informed neural networks serve as data-efficient surrogates for complex dynamical systems. Beyond predictive accuracy, the proposed NASP PINN framework provides both a computational solver and an interpretive model, underscoring its dual role as a practical simulator and a theoretical tool for analyzing noise-driven dynamics. It thus reaffirms its value as a powerful and general approach for studying complex stochastic phenomena in which purely analytical or numerical methods may be insufficient.
@@ Acknowledgements
$\alpha$ Learning rate (used in $\text{Adam}(\theta,\alpha)$)
@@
$\theta$ Model parameters
@@ 16 $\tfrac{1}{2}\,\sigma^{2}\log(\varepsilon^{-1})\leftarrow\text{Deterministic timescale}$ ;
The codes developed and used to generate the data supporting the findings of this study are publicly accessible here [58].

### v2

@@ 5.4 Model, optimizer setup, and hyperparameters
Training is performed via backpropagation using the Adam optimizer, see the algorithm in the Appendix Appendix. We initialize a StatePredictor with three 128-unit Tanh blocks and a dynamics head and set $dt=0.05$, epochs $1\times 10^{4}$, batch size =512=512, and learning rate $1.0\times 10^{-3}$. The total loss is a weighted sum of the data loss, the initial-condition loss, the residual loss, and the barrier-based physics loss weighted by, $\lambda_{\tiny{data}},\lambda_{\tiny{ic}},\lambda_{\tiny{phy1}},\lambda_{\tiny{phy2}}$ respectively. During training, each minibatch is a consecutive time sequence. Because the loss weights $\lambda_{i}$ are dynamically adjusted during training, we additionally report how much each loss component contributes to the parameter updates over time. For each component $\pazocal{L}_{i}(\theta)$, we compute the gradient-weighted share
$s_{i}(\text{epoch})\;=\;\frac{\lambda_{i}\,\lVert\nabla_{\theta}\pazocal{L}_{i}\rVert}{\sum_{j}\lambda_{j}\,\lVert\nabla_{\theta}\pazocal{L}_{j}\rVert},$ (22)
so that $\sum_{i}s_{i}=1$ at every epoch, where $i,j\in\{\text{data, ic, phy1, phy2}\}$. This diagnostic reveals which term dominates the optimization at different phases of learning [61].
For evaluation, we predict the target sequence and compute the normalized root-mean-square error (NRMSE). The optimal training epoch is selected as the one yielding the smallest NRMSE, defined for a reference sequence $y=\{y_{t}\}_{t=1}^{N}$ and its predictions $\hat{y}=\{\hat{y}_{t}\}_{t=1}^{N}$ by
@@
Here $y=(v(t),w(t))$ denotes the two-dimensional reference trajectory of the system state, and $N$ is the number of time points in the evaluated prediction window. The NRMSE therefore aggregates errors across both the fast variable $v$ and the slow variable $w$, reflecting the need for the surrogate to jointly predict both components of the slow–fast dynamics. Lower NRMSE values indicate better predictive performance; consequently, a faster decrease in NRMSE during training corresponds to faster convergence of the optimization process.
It is worth mentioning that several training experiments (not shown) were performed using NumPy and PyTorch with different initialization seeds and all showed the same qualitative convergence and rollout behavior, confirming that the NASP–PINN is not sensitive to different initializations.
@@ 5.5 Results on training and evaluation
We evaluate the NASP–PINN on held-out sequences to measure both one-step prediction quality and multi-step rollout fidelity. In addition, we conduct a formal ablation study to quantify the individual contribution of each physics-informed component of the loss. Specifically, we benchmark four model variants: (i) data only, (ii) data + ic + phy1 (SDE residual loss), (iii) data + phy2 (Kramers barrier-matching), and (iv) the full model with all terms included. This ablation design isolates the effect of each physics term on optimization behavior, predictive accuracy, and out-of-sample performance.
For each experiment, the stochastic FHN system is integrated over a fixed time horizon $t\in[0,25\times 10^{3}]$, which produces $5\times 10^{5}$ data points with $\Delta t=0.05$. The training window corresponds to the initial segment $t\in[0,10\times 10^{3}]$, processed in consecutive minibatches as described in Section 5.4. The held-out test sequence refers to the immediately following portion of the same simulated trajectory, i.e., $t\in(10\times 10^{3},25\times 10^{3}]$, which is not used during training and therefore serves as unseen future dynamics for evaluating temporal generalization. The lengths of the training and testing windows used in all quantitative evaluations are identical.
To compute the NRMSE (Eq. (23)), a ground-truth reference trajectory of the same length as the model prediction is required. We therefore designate the trajectory obtained by numerically integrating Eq. (1) as the reference sequence, and evaluate all model variants against this same trajectory. All model variants are evaluated against this same reference, ensuring a consistent and fair comparison across ablation settings. This reference sequence is representative because it is drawn from the stationary stochastic dynamics, and SISR statistics (such as the ISI distribution and the slow drift of $w$) remain stable for fixed parameter values used.
The results of this training and evaluation procedure, including the ablation study, are summarized in Fig. 8 and Table 1. Table 1 reports the train and test NRMSE for the training window $t\in[0,10\times 10^{3}]$ and the held-out test window $t\in(10\times 10^{3},25\times 10^{3}]$, using the same reference sequence and evaluation protocol for all ablation variants. Figure 9 illustrates a representative free-rollout under this same train/test setup, providing a qualitative view of long-horizon behavior—including spike timing, phase stability, and amplitude preservation—that complements the quantitative performance reported in Table 1.
Figure 8(a) presents the results of the ablation study using the training NRMSE, computed over both state variables $(v,w)$ to capture errors in both the fast spiking dynamics and the slow recovery drift. We report the NRMSE during training because it provides a physically interpretable and comparable measure of prediction quality across the different loss configurations, thereby highlighting differences in optimization speed and stability. The four curves correspond to the data-only model (red), data+ic+phy1 (yellow), data+phy2 (blue), and the full data+ic+phy1+phy2 model (green). Incorporating either of the physics-informed terms (phy1 or phy2) accelerates convergence and lowers the asymptotic error relative to the purely data-driven baseline. The full NASP–PINN achieves the fastest and most stable convergence as well as the lowest NRMSE, consistent with the quantitative train/test results reported in Table 1.
Figure 8(b) illustrates how the dynamically weighted objective is optimized over time. Early epochs are dominated by the supervised data and initial-condition terms, reflecting rapid alignment to observed one-step targets. As training progresses, the optimization emphasis shifts toward the physics constraints, with phy1 and phy2 accounting for the majority of the gradient budget in later epochs. This indicates that the physics terms are not negligible and they actively shape the learned dynamics by enforcing local SDE consistency (phy1) and global timescale structure (phy2), which is consistent with the lower test NRMSE of the full model reported in Table 1.
@@
Figure 8: (a) NRMSE values during training. Ablation study of NASP–PINN training performance under different loss configurations. Shown are NRMSE curves for four variants: data only (red), data+ic+phy1 (yellow), data+phy2 (blue), and data+ic+phy1+phy2 (green). The physics components (phy1, phy2) significantly accelerate convergence and reduce error. The full model exhibits the best optimization behavior, achieving the fastest convergence, most stable training, and the best test performance (see Table 1). (b) Effective gradient contribution of each loss component for the full model, computed as $s_{i}$ (Eq. (22)). Shows how the optimization focus transitions across training, indicating which constraints dominate parameter updates over time.
@@
Table 1 summarizes train/test NRMSE for the same variants. Data-only training yields the weakest train/test performance (0.053 train, 0.062 test), indicating mild overfitting to short-horizon local patterns. Enforcing dynamical residuals (phy1) and/or barrier-matching (phy2) improves test performance and lowers both errors (0.040–0.042 train, 0.050 test). The full loss achieves the best results by a clear margin (0.027 train, 0.035 test), demonstrating that the two physics priors are complementary: phy1 aligns local derivatives with the SDE, while phy2 shapes global exit statistics via timescale matching, together constraining the hypothesis class to trajectories that are both locally consistent and globally coherent. Relative to data-only training, the PINN reduces test NRMSE by approximately 43% (0.062 $\rightarrow$ 0.035) while also quickly lowering train error, consistent with a better-conditioned objective and stronger inductive bias.
@@
Table 1: NRMSE for NASP variants under different loss combinations.
@@
From a learning-theoretic perspective, the physics terms act as inductive biases and implicit regularizers. phy1 reduces variance by penalizing derivative mismatches estimated via automatic differentiation; phy2 reduces bias in the noise-triggered event timing by anchoring the learned dynamics to the Arrhenius-type scaling required for SISR (timescale matching). Their combination explains the lower test error and the smoother, faster descent curves in Fig. 8(a). In broader machine-learning terms, these physics-based penalties constrain the hypothesis space in a way analogous to structural regularization in deep sequence models, improving both data efficiency and temporal consistency.
To test stability beyond the training horizon, we perform open-loop (free-rollout) simulations and compare against direct stochastic simulations. Figure 9 compares the time series of $v(t)$ and $w(t)$ for the data-only model (panels (a1)–(a2)) and the full NASP–PINN (panels (b1)–(b2)) across both the training and prediction windows. Over the horizon shown, the two models can appear visually similar, as both track the large-amplitude spikes in $v$ and the slow drift in $w$. However, the full NASP–PINN achieves a substantially lower test NRMSE ($0.035$ vs. $0.062$ for data-only, see Table 1), representing a $43\%$ improvement in predictive accuracy. In slow–fast excitable systems, such differences can be highly consequential: even small local errors accumulate along the slow manifold and lead to progressively shifting spike times during very long rollouts. Because SISR coherence depends sensitively on the regularity and alignment of these escape events (precision of the spiking times), the physics-informed model would definitively yield markedly more reliable very long-horizon dynamics.
The qualitative agreement in panels (b1)–(b2) demonstrates that the learned integrator is both noise-aware and stiffness-compatible, properties that are crucial for accurately resolving slow–fast SDEs. The barrier-matching term (phy2) teaches the model where and at what rate escapes occur in $w$, whereas the residual term (phy1) enforces local dynamical consistency between successive steps. Together, these constraints produce a surrogate that maintains phase alignment and amplitude fidelity far more robustly than the data-only model, despite the apparent visual similarity over relatively short windows.
Figure 9: Comparison of rollout performance for the data-only model (top row) and the full NASP–PINN (bottom row). Panels (a1)–(a2) show the predicted and true time series of $v(t)$ and $w(t)$ for the data-only model. Panels (b1)–(b2) show the corresponding results for the physics-informed model. Although the trajectories appear visually similar over short horizons, the NASP–PINN achieves a substantially lower test NRMSE ($0.035$ vs. $0.062$), a $43\%$ improvement. In slow–fast excitable systems, such errors accumulate over time and lead to shifts in spike timing that directly affect SISR coherence.
@@
Overall, the findings demonstrate that embedding the governing stochastic dynamics and the SISR-specific constraints directly into the loss function transforms NASP into a data-efficient and generalizable framework for modeling noise-driven multiscale dynamics. Across all evaluation metrics—including learning curves, final NRMSE, and long-horizon rollouts—the Data + ic + phy1 + phy2 configuration consistently yields superior performance, with the most pronounced improvements observed in extended predictions where preserving the intrinsic slow–fast structure and accurate escape timing is critical for reproducing SISR. Physics-informed training, particularly the combination of local residual constraints and global barrier-matching, thus serves as an effective inductive bias that enhances generalization while ensuring the model faithfully respects the physics that drives self-induced stochastic resonance.
@@
Figure 10(a)-(b) compares the noise–coherence relationship predicted by the NASP-PINN with that obtained from direct stochastic simulations of the FHN model Eq. (1). The $\mathrm{CV}$ of interspike intervals is plotted as a function of noise intensity $\sigma$ for different values of the excitability parameter $a$ and the timescale separation $\varepsilon$. Both panels show that the NASP-PINN does well in predicting the degree of coherence due SISR, as $\sigma$, $a$, and $\varepsilon$ are varied.
The close quantitative agreement between both sets of curves demonstrates that the PINN accurately learns the underlying stochastic dynamics and correctly captures how the coherence of SISR depends on excitability and timescale separation. In particular, the network reproduces the characteristic shift of the $\mathrm{CV}$ minima toward higher noise intensities as $a$ or $\varepsilon$ increase, reflecting the loss of timescale matching between deterministic relaxation and stochastic escape predicted by the theoretical SISR condition. This confirms that embedding the governing physics directly into the training objective enables the PINN to generalize beyond the training data and faithfully recover the resonance structure of the stochastic FHN system.
Figure 10: True ($\bullet$ marker) and the NASP-PINN-predicted ($\star$ marker) coherence degree of SISR. (a) and (b) show $\mathrm{CV}$ versus noise intensity $\sigma$ for varying excitability $a$ and timescale separation $\varepsilon$ parameters, respectively, demonstrating the ability of the NASP-PINN to predict with a high degree of accuracy the dependence of SISR on both parameters. Other parameters: $b=1.0$, $c=2.0$.
@@ 6 Summary and conclusions
Building on these theoretical insights, we introduced a Physics-Informed Neural Network (PINN) based on a Noise-Augmented State Predictor (NASP) architecture to learn the stochastic FHN dynamics directly from data. The composite loss integrates data fidelity, dynamical residuals, and the physics-based constraints derived from SISR theory. Through a dedicated ablation study, we quantified the individual contributions of each loss component—data term, SDE residual (phy1), and Kramers timescale-matching constraint (phy2)—and observed that each provides a meaningful inductive bias. In particular, combining both physics-informed terms yielded the fastest convergence, lowest error, and strongest predictive stability compared with purely data-driven approaches. The trained NASP–PINN accurately reproduced long-term SISR dynamics and captured the dependence of the $\mathrm{CV}$ on noise intensity, excitability, and timescale separation, demonstrating its ability to faithfully encode the underlying stochastic physics.
Although in this work the governing SDEs of the FHN neuron are fully known and therefore directly incorporated as physical priors (phy1 and phy2), the framework extends naturally to scenarios where the equations are partially known or only qualitatively specified. In such cases, data-only terms can be used to fit the unknown components, while the available physics—such as conservation laws, timescale separation, escape-rate scaling, or qualitative dynamical structure—can still be incorporated as weak constraints or surrogate regularizers. Thus, the proposed NASP–PINN model is not restricted to analytically tractable systems, but can act as a hybrid data–physics learner for experimental recordings where only partial mechanistic knowledge is available.
@@
In addition, this physics-informed neural framework aligns with current trends in scientific machine learning, where neural ordinary differential equations (neural ODEs) [59], reservoir computing [60], and physics-informed neural networks serve as data-efficient surrogates for complex dynamical systems. Beyond predictive accuracy, the proposed NASP-PINN framework provides both a computational solver and an interpretive model, underscoring its dual role as a practical simulator and a theoretical tool for analyzing noise-driven dynamics. It thus reaffirms its value as a powerful and general approach for studying complex stochastic phenomena in which purely analytical or numerical methods may be insufficient.
Future work will explore the unknown (or partially known) model setting, in which neural networks learn dynamics using drifts or noise terms identified through sparse regression methods (e.g., SINDy), and physics-based SISR constraints regularize the solution space. This direction will enable application of the framework to biological neuron data, where no (or only partial) mathematical models are available.
@@ Acknowledgements
$\theta$ Model parameters (weights and biases of the MLP)
$\alpha$ Learning rate (used in $\text{Adam}(\theta,\alpha)$)
@@ 16 $\tfrac{1}{2}\,\sigma^{2}\log(\varepsilon^{-1})\leftarrow\text{Deterministic timescale}$ ;
The codes developed and used to generate the data supporting the findings of this study are publicly accessible [62].